OpenStudy (anonymous):

What is the slope of the line that passes through the pair points? (-6, 8), (2, 3) -3/8 -1/8 -7/8 -5/8

OpenStudy (the_fizicx99):

\[M = y2-y1/x2-x1\]

OpenStudy (anonymous):

So it would be (3 - 8) / (2 - (-6))

OpenStudy (the_fizicx99):

Yes, - and - = + so??

OpenStudy (anonymous):

-5 / 8 ...?

OpenStudy (the_fizicx99):

Yesh

OpenStudy (the_fizicx99):

Or, -0.625

OpenStudy (anonymous):

-0.625... How would we put that in a fraction?

OpenStudy (the_fizicx99):

-0.625/1

OpenStudy (anonymous):

-5/8 ?

OpenStudy (the_fizicx99):

-5/8 = -0.625 -0.625 as a fraction would be -0.625/1

OpenStudy (anonymous):

Yeah it's just cuz my only options are -3/8 -1/8 -7/8 -5/8

OpenStudy (the_fizicx99):

Then D,

OpenStudy (anonymous):

Thank you so much ^-^

OpenStudy (the_fizicx99):

Yw :)